An air conditioning contractor has kept detailed records on the lifetimes of a random sample of 250 compressors. She plans to use the sample mean lifetime, x, of those 250 compressors as her estimate or he population mea lifetime μ o al such compressors if the e meso this randon compressor avea standard deviation of 30 months, what is the probability that the contractors estimate will be within 7 months of the true mean? The probability is (Round to four decimal places as needed.)
Answer #1

Sample size = n = 250

Mean = mu = 250

Standard deviation = sigma = 30

We have to find the probability that the contractor's estimate will be within 7 months of the true mean.

That is we have to find P( 250 - 7 < x < 250 + 7) = P(243 < x < 257)

For finding this probability we have to find z score.

243-250 30 x-μ 0.23

257-250 30 x-μ 0.23

That is we have to find P( - 0.23 < Z < 0.23)

P( - 0.23 < Z < 0.23) = P(Z < 0.23) - P(Z < - 0.23) = 0.5910 - 0.4078 = 0.1832